You drove 365 miles using 15 gallons of gasoline a that costs $3.50 per gallon. If you get 27 mi/ gal on the highway  and 19 in the city, how much did you spend for fuel for highway driving and how much for city driving? 
Answer by DrBeeee(684)

I had this all worked out for you and the puter burped and I lost the detailed answer. Here it is quickly w/o detail.
Let c = gallons used in city
Let h = gallons used in highway
(1) 19c + 27h = 365
(2) c + h = 15
Solution of (1) and (2)
c = 5 and h = 10
Cost in city is $17.50
Cost on highway is $35.00